A production facility contains two machines that are used to rework items that are initially defective. Let 𝑋 be the number of hours that the first machine is in use and let π‘Œ be the number of hours that the second machine is in use, on a randomly chosen day. Assume that 𝑋 and π‘Œ have a joint probability density function given by 𝑓(π‘₯) = { 3 2 (π‘₯ 2 + 𝑦 2 ) 0 < π‘₯ < 1 π‘Žπ‘›π‘‘ 0 < 𝑦 < 1 0 π‘œπ‘‘β„Žπ‘’π‘Ÿπ‘€π‘–π‘ π‘’. a. What is the probability that both machines are in operation for less than half an hour?

1
Expert's answer
2022-01-28T09:24:33-0500
"P(X<0.5, Y<0.5)=\displaystyle\int_{0}^{0.5}\displaystyle\int_{0}^{0.5}\dfrac{3}{2}(x^2+y^2)dydx""=\dfrac{3}{2}\displaystyle\int_{0}^{0.5}[x^2y+\dfrac{y^3}{3}]dx\begin{matrix}
 0.5 \\
 0
\end{matrix}""=\dfrac{3}{2}\displaystyle\int_{0}^{0.5}(\dfrac{1}{2}x^2+\dfrac{1}{24})dx""=\dfrac{3}{2}[\dfrac{1}{6}x^3+\dfrac{1}{24}x]\begin{matrix}
 0.5 \\
 0
\end{matrix}""=\dfrac{3}{2}(\dfrac{1}{48}+\dfrac{1}{48})=\dfrac{1}{16}""P(X<0.5, Y<0.5)=\dfrac{1}{16}"